OpenStudy (anonymous):

Given: f(x) = 2x – 5 and g(x) = x + 2, find (f º g)(x) Select one: a. 4x – 3 b. 2x – 6 c. 2x – 14 d. 2x – 1

OpenStudy (anonymous):

man we have dexter and batman in here

OpenStudy (anonymous):

you dare mock the dark night?

OpenStudy (anonymous):

Sorry it's the same thing as f(g(x)) \[f(g(x))= 2(x+2)-5\]

OpenStudy (anonymous):

I was thinking of something else before xD

OpenStudy (anonymous):

so its d

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

thanks i didnt quite get that one

OpenStudy (anonymous):

This is an composite function \[(f\cdot g)(x)=f[g(x)]\]What does it means?!\[f(x)=2x-5\]Now we have to replace x by g(x)\[f[g(x)]=2*g(x)-5\]Replecing the function g(x)\[f(x+2)=2*(x+2)-5\]\[f(x+2)=2x+4-5\]the answer is\[\boxed{(f\cdot g)(x)=2x-1}\]